# Locally homogeneous Axiom A flows II: geometric structures for Anosov subgroups

### v2

@@ 7.1. Invariant subsets of the flow space
As before, let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, with transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, and let $\Gamma<G$ be $\Theta$-Anosov. Consider a flow space ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ as defined in Section 6. Then it turns out that $\Gamma$ possesses a non-empty flow-invariant domain of proper discontinuity in ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ for suitable parameters $\mathbf{s}\in\mathbb{R}^{\Theta}$. The heart of the construction is the following
@@
~$\widetilde{\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}}$ $:=p^{-1}(\Omega_{\Gamma})=\left\{\mu\in\mathbb{L}^{\mathbf{s}}_{\Theta}\,\middle|\,p(\mu)\in\Omega_{\Gamma}\right\},$
@@
Note that ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ (resp. $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$) is the restriction of the bundle $\mathbb{L}^{\mathbf{s}}_{\Theta}$ to the subset $\Omega_{\Gamma}$ (resp. $\Lambda^{\pitchfork}_{\Gamma}$) of its base space ${{\mathcal{F}}^{\pitchfork}}$. In particular, since the flow $\phi^{t}$ preserves the fibers of $\mathbb{L}^{\mathbf{s}}_{\Theta}$, we see that ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ are $\phi^{t}$-invariant subsets of $\mathbb{L}^{\mathbf{s}}_{\Theta}$. Further elementary properties can be shown exactly as in the projective Anosov case, see [DMS24, Lemmas 3.4, 3.7]:
@@
The sets $\Omega_{\Gamma}\subset{{\mathcal{F}}^{\pitchfork}}$ and ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ are open and $\Gamma$-invariant. The set $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ is closed, $\Gamma$-invariant, and contained in ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
In particular, since $\Lambda^{\pitchfork}_{\Gamma}$, and consequently $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$, is non-empty unless $\Gamma$ is finite, in which case $\Omega_{\Gamma}$ is all of ${{\mathcal{F}}^{\pitchfork}}$, we conclude that the sets $\Omega_{\Gamma}$ and ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ are always non-empty.
@@ 7.2. Admissible parameters
[…]
@@ 7.4. Proof of proper discontinuity
Let $\Gamma<G$ be $\Theta$-Anosov and consider a $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}\in\mathbb{R}^{\Theta}$. Then $\Gamma$ acts properly discontinuously on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and cocompactly on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$. If $\Gamma$ is torsion-free, then the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is free.
@@
Theorem 1 is a statement about ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ as $\Gamma$-spaces, it does not involve the flow. In particular, (6.4) implies that
[…]
@@
for any $\gamma\in\Gamma$ with $\lambda(\gamma)\neq 0$ and any norm on ${\bf\mathfrak{a}}$, the compactness of $\left\{X\in\mathcal{L}(\Gamma)\,\middle|\,\left\|X\right\|=1\right\}$ and the $(\Gamma,\Theta)$-admissibility of $\mathbf{s}$ mean that it suffices to show that $\left\|\lambda(\gamma_{k})\right\|\to\infty$, i.e., $\lambda(\gamma_{k})\to\infty$ in ${\bf\mathfrak{a}}$, for any sequence $(\gamma_{k})$ such as in the statement. To this end, we first apply Lemma 2.5.(2), which says that the stable length $|\gamma_{k}|_{\infty}$ converges to $\infty$. Then by Definition 7.1 we have that $\alpha(\lambda(\gamma_{k}))\to+\infty$, in particular $\lambda(\gamma_{k})\to\infty$ in ${\bf\mathfrak{a}}$. ∎
@@
$\mathcal{C}_{\sigma^{\prime}}(\mathbf{y},g)=\sum_{\alpha\in\Theta}\mathcal{C}_{\sigma^{\alpha}}(\Pi^{{\mathcal{F}}^{-}}_{\alpha}(\mathbf{y}),g)\quad\forall\,(\mathbf{y},g)\in{\mathcal{F}}^{-}\times G.\vspace*{-1em}$
@@
The cocompactness of the action on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ comes from Proposition 7.7, and any properly discontinuous action of a torsion-free group is free, so all that is left to prove is the proper discontinuity of the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
@@
Now, extract a further subsequence so that $\gamma_{k}$ converges to $\gamma_{+}\in\partial_{\infty}\Gamma$ and $\gamma_{k}^{-1}$ converges to $\gamma_{-}\in\partial_{\infty}\Gamma$ (note that $\gamma_{+}\neq\gamma_{-}$), and denote by $(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}$ the base point of $(\nu^{+},\nu^{-})$. By virtue of $(\nu^{+},\nu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$, either $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ or $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$. The first case $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ implies via Lemma 7.11 and Proposition 7.8 that $\gamma_{k}\cdot\nu^{-}_{k}\rightarrow\infty$ in ${\mathscr{D}}^{\mathbf{s}}_{>0}{\mathcal{F}}^{-}_{\bullet}$, which is absurd because of the assumption $\gamma_{k}\cdot(\nu^{+}_{k},\nu^{-}_{k})\rightarrow(\mu^{+},\mu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. In the second case $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$ we can repeat the same argument with $(\Theta,\mathbf{s})$ replaced by $(\iota_{o}(\Theta),\iota(\mathbf{s}))$, as is justified by (7.2), (7.5), and (7.9). This leads to the analogous contradiction, and therefore the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is properly discontinuous. ∎
@@
Note that our proof of Theorem 1 never explicitly uses the coarsely linear growth defining the $\Theta$-Anosov property in Definition 7.1: the only step in which we use the growth of $\mathcal{P}_{\alpha}(\gamma_{k})$ is at the end of the proof of Proposition 7.8, where we only need that $\mathcal{P}_{\alpha}(\gamma_{k})\to+\infty$ as $k\to\infty.$ Subgroups $\Gamma<G$ satisfying this property are called $\Theta$-transverse, and were discussed in the introduction. In the proof of Lemma 7.9 we use in addition the existence of a pair of continuous, transverse, equivariant, dynamics-preserving boundary maps, and in total this implies the $\Theta$-Anosov property [GGKW17, Thm. 1.7].
@@ 8. Dynamics on the quotient manifold
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ the transverse flag space, $\Gamma<G$ a torsion-free $\Theta$-Anosov subgroup, and $\mathbf{s}\in\mathbb{R}^{\Theta}$ a $(\Gamma,\Theta)$-admissible parameter. Then we know from Theorem 1 that $\Gamma$ acts properly discontinuously and freely on the $\phi^{t}$-invariant open set ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ and cocompactly on the $\phi^{t}$-invariant closed set $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. Since the $\Gamma$-actions on these sets commute with the flow $\phi^{t}$, the quotient manifold
@@ 8.1. Simplified notations
and their projections on the base space ${{\mathcal{F}}^{\pitchfork}}$ will be written as
$\big{(}[\mu^{+}],[\mu^{-}]\big{)}\in{{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\subset{\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\,.$
@@ 8.2. Topological dynamics
Consider sequences $x_{k}\in{\mathcal{M}}$ and $t_{k}\to+\infty$ and points $x,y\in{\mathcal{M}}$ such that $x_{k}\to x\in{\mathcal{M}}$ and $\phi^{t_{k}}(x_{k})\to y\in{\mathcal{M}}$. Let $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ (resp. $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$) be a lift of $x$ (resp. of $y\in{\mathcal{M}}$). Then $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $[\mu^{-}]\in\Lambda^{-}_{\Gamma}$.
@@
Assuming that $\xi^{+}(\gamma_{-})$ is transverse to $[\mu^{-}]$ leads to a contradiction: since $\mu_{k}^{-}\to\mu^{-}\neq 0$, Lemma 7.11 would imply that $\gamma_{k}\mu_{k}^{-}\to 0$, yet $\gamma_{k}\mu_{k}^{-}=e^{t_{k}}f\nu_{k}^{-}$ and $\nu_{k}^{-}\to\nu^{-}\neq 0$. Since $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$, we must have $[\mu^{+}]\pitchfork\xi^{-}(\gamma_{-})$, therefore $\gamma_{k}[\mu^{+}_{k}]\to\xi^{+}(\gamma_{+})$ by (A3). This means that $f[\nu^{+}]=\xi^{+}(\gamma_{+})$, and in particular $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$.
This implies that $\xi^{+}(\gamma_{+})\pitchfork f[\nu^{-}]$, so $\gamma_{k}^{-1}f[\nu^{-}_{k}]\to\xi^{-}(\gamma_{-})$, i.e. $[\mu^{-}]=\xi^{-}(\gamma_{-})\in\Lambda^{-}_{\Gamma}$. ∎
@@
Let $x\in{\mathcal{M}}$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Then $x\in\mathrm{Per}(\phi^{t})$ if and only if there is $\gamma\in\Gamma\setminus\{e\}$ such that $[\nu^{+}]=\xi^{+}(\gamma^{+})$ and $[\nu^{-}]=\xi^{-}(\gamma^{-})$. [^3] Recall that we assume $\Gamma$ to be torsion-free here, so that any $\gamma\in\Gamma\setminus\{e\}$ has infinite order. In this case, if $\gamma$ is primitive, then the period of $x$ is $\mathcal{J}^{\mathbf{s}}_{\Theta}(\gamma)$.
@@
Assume that $x\in\mathrm{Per}(\phi^{t})$, and let $T>0$ be such that $\phi^{T}(x)=x$. There is $\gamma\in\Gamma\setminus\{e\}$ such that $(e^{-T}\nu^{+},e^{T}\nu^{-})=\gamma\cdot(\nu^{+},\nu^{-})$. If $[\nu^{-}]$ were to be transverse to $\xi^{+}(\gamma^{-})$, then Lemma 7.11 would imply that $\gamma^{k}\cdot\nu^{-}\to 0$ as $k\to+\infty$, which is a contradiction with $\gamma^{k}\cdot\nu^{-}=e^{kT}\nu^{-}$. As $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$, we must have $[\nu^{+}]\pitchfork\xi^{-}(\gamma^{-})$, so $\gamma^{k}[\nu^{+}]\to\xi^{+}(\gamma^{+})$ as $k\to+\infty$, i.e. $[\nu^{+}]=\xi^{+}(\gamma^{+})$. Since $[\nu^{-}]\pitchfork[\nu^{+}]=\xi^{+}(\gamma^{+})$, we also find that $\gamma^{k}[\nu^{-}]\to\xi^{-}(\gamma^{-})$ as $k\to-\infty$, therefore $[\nu^{-}]=\xi^{-}(\gamma^{-})$.
If the assumption is now that $[\nu^{+}]=\xi^{+}(\gamma^{+})$ and $[\nu^{-}]=\xi^{-}(\gamma^{-})$ for some element $\gamma\in\Gamma\setminus\{e\}$, then combining (6.5), (3.21) and (3.19) yields $\phi^{w_{{\mathbf{s}},\Theta}(\lambda(\gamma))}(x)=x$. ∎
@@
Any flow satisfies $\overline{\mathrm{Per}(\phi^{t})}\subset\mathcal{NW}(\phi^{t})$. Let $x\in\mathcal{NW}(\phi^{t})$, and consider sequences $x_{k}\to x$ in ${\mathcal{M}}$ and $t_{k}\to+\infty$ in $\mathbb{R}$ such that $\phi^{t_{k}}(x_{k})\to x$. If $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ is a lift of $x$, then Lemma 8.1 shows that $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $[\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$, i.e. $\mathcal{NW}(\phi^{t})\subset\mathcal{K}$.
@@
Let $U\subset{\mathcal{M}}$ be a relatively compact subset containing $\mathcal{K}$, and consider $y\in\overline{\bigcap_{t\in\mathbb{R}}\phi^{t}(U)}$ and a lift $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$. Consider first sequences $t_{k}\to+\infty$ and $x_{k}\in U$ such that $\phi^{t_{k}}(x_{k})\to y$, and assume without loss of generality that $x_{k}$ has a limit $x\in{\mathcal{M}}$. Lemma 8.1 shows that $[\mu^{-}]\in\Lambda^{-}_{\Gamma}$. Consider now sequences $s_{k}\to-\infty$ and $z_{k}\in U$ such that $y=\lim\phi^{s_{k}}(z_{k})$, and assume that $z_{k}\to z\in{\mathcal{M}}$. Applying Lemma 8.1 to $\phi^{-s_{k}}\left(\phi^{s_{k}}(y)\right)$ shows that $[\mu^{+}]\in\Lambda^{+}_{\Gamma}$, hence $y\in\mathcal{K}$.
@@
Since $\mathcal{K}$ is compact and $\phi^{t}$-invariant, we automatically get $\mathcal{K}\subset{\mathcal{T}}(\phi^{t})$. Let $x\in{\mathcal{T}}(\phi^{t})$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Let $y\in{\mathcal{M}}$ be an $\omega$-limit point of $x$, i.e. $y=\lim_{k\to+\infty}\phi^{t_{k}}(x)$ for some sequence $t_{k}\to+\infty$ (it exists because $x$ is trapped). Then Lemma 8.1 shows that $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$. The same reasoning applied to an $\alpha$-limit point of $x$ (i.e. some $\lim\phi^{s_{k}}(x)$ with $s_{k}\to-\infty$) shows that $[\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$.
It remains to prove that the restriction of the flow $\phi^{t}$ to $\mathcal{K}$ is topologically transitive. This is again done by complete analogy to the projective case [DMS24, Lemma 3.12]: By [Gro87, 8.2.H], we may consider $(a,b)\in\partial_{\infty}\Gamma^{(2)}$ whose $\Gamma$-orbit is dense in $\partial_{\infty}\Gamma^{(2)}$. Let $x\in\mathcal{K}$ be an element admitting a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{K}}$ with $[\nu^{+}]=\xi(a)$ and $[\nu^{-}]=\xi^{*}(b)$. Let $y\in\mathcal{K}$ and consider a lift $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{K}}$. Let $U\subset\mathcal{K}$ be an open subset containing $y$, and $\widetilde{U}\subset\widetilde{\mathcal{K}}$ its preimage. As the bundle projection $p:{\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$ is an open map, there is $\gamma\in\Gamma$ such that $\gamma\cdot([\mu^{+}],[\mu^{-}])\in p(\widetilde{U})$, i.e., there is $t\in\mathbb{R}$ such that $\gamma\cdot(e^{-t}\mu^{+},e^{t}\mu^{-})\in\widetilde{U}$, hence $\phi^{t}(x)\in U$. ∎

### v3

@@ 7.1. Invariant subsets of the flow space
[…]
@@
$\widetilde{\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}}$ $:=p^{-1}(\Omega_{\Gamma})=\left\{\mu\in\mathbb{L}^{\mathbf{s}}_{\Theta}\,\middle|\,p(\mu)\in\Omega_{\Gamma}\right\},$
@@
[…]
@@
The sets $\Omega_{\Gamma}\subset{{\mathcal{F}}^{\pitchfork}}$ and $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ are open and $\Gamma$-invariant. The set $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ is closed, $\Gamma$-invariant, and contained in $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
[…]
@@ 7.2. Admissible parameters
$\mathbf{s}\text{ is }(\Gamma,\Theta)\text{-admissible}\iff\iota(\mathbf{s})\text{ is }(\Gamma,\iota_{o}(\Theta))\text{-admissible}.$ (7.5)
@@ 7.4. Proof of proper discontinuity
Let $\Gamma<G$ be $\Theta$-Anosov and consider a $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}\in\mathbb{R}^{\Theta}$. Then $\Gamma$ acts properly discontinuously on $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and cocompactly on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$. If $\Gamma$ is torsion-free, then the $\Gamma$-action on $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is free.
@@
Theorem 1 is a statement about $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ as $\Gamma$-spaces, it does not involve the flow. In particular, (6.4) implies that
$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\simeq\widetilde{{\mathcal{M}}}_{\Gamma,\iota_{o}(\Theta)}^{\iota(\mathbf{s})}\quad\text{as $\Gamma$-manifolds.}$ (7.9)
@@
for any $\gamma\in\Gamma$ with $\lambda(\gamma)\neq 0$ and any norm on $\bf\mathfrak{a}$}, the compactness of $\left\{X\in\mathcal{L}(\Gamma)\,\middle|\,\left\|X\right\|=1\right\}$ and the $(\Gamma,\Theta)$-admissibility of $\mathbf{s}$ mean that it suffices to show that $\left\|\lambda(\gamma_{k})\right\|\to\infty$, i.e., $\lambda(\gamma_{k})\to\infty$ in $\bf\mathfrak{a}$}, for any sequence $(\gamma_{k})$ such as in the statement. This follows from Lemma 7.2. ∎
@@
$\mathcal{C}_{\sigma^{\prime}}(\mathbf{y},g)=\sum_{\alpha\in\Theta}\mathcal{C}_{\sigma^{\alpha}}(\Pi^{{\mathcal{F}}^{-}}_{\alpha}(\mathbf{y}),g)\quad\forall\,(\mathbf{y},g)\in{\mathcal{F}}^{-}\times G.\vskip-10.00002pt$
@@
The cocompactness of the action on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ comes from Proposition 7.8, and any properly discontinuous action of a torsion-free group is free, so all that is left to prove is the proper discontinuity of the $\Gamma$-action on $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
@@
Now, extract a further subsequence so that $\gamma_{k}$ converges to $\gamma_{+}\in\partial_{\infty}\Gamma$ and $\gamma_{k}^{-1}$ converges to $\gamma_{-}\in\partial_{\infty}\Gamma$ (note that $\gamma_{+}\neq\gamma_{-}$), and denote by $(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}$ the base point of $(\nu^{+},\nu^{-})$. By virtue of $(\nu^{+},\nu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$, either $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ or $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$. The first case $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ implies via Lemma 7.12 and Proposition 7.9 that $\gamma_{k}\cdot\nu^{-}_{k}\rightarrow\infty$ in ${\mathscr{D}}^{\mathbf{s}}_{>0}{\mathcal{F}}^{-}_{\bullet}$, which is absurd because of the assumption $\gamma_{k}\cdot(\nu^{+}_{k},\nu^{-}_{k})\rightarrow(\mu^{+},\mu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. In the second case $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$ we can repeat the same argument with $(\Theta,\mathbf{s})$ replaced by $(\iota_{o}(\Theta),\iota(\mathbf{s}))$, as is justified by (7.2), (7.5), and (7.9). This leads to the analogous contradiction, and therefore the $\Gamma$-action on $\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is properly discontinuous. ∎
@@
We used properties of $\Theta$-Anosov subgroups in two different ways when proving Theorem 1: to give a definition of the proper discontinuity domain (Definition 7.3 involves the limit maps) and to prove proper discontinuity. We expect that Theorem 1 is still valid for a larger class of discrete subgroups, provided they have a good notion of limit set in $\mathcal{F}^{\pm}$ and growth properties of Jordan projections such as in Lemma 7.2, as is the case for $\Theta$-transverse subgroups discussed in the introduction.
@@ 8. Dynamics on the quotient manifold
[…]
@@ 8.1. Simplified notations
[…]
@@ 8.2. Topological dynamics
Consider sequences $x_{k}\in{\mathcal{M}}$ and $t_{k}\to+\infty$ and points $x,y\in{\mathcal{M}}$ such that $x_{k}\to x\in{\mathcal{M}}$ and $\phi^{t_{k}}(x_{k})\to y\in{\mathcal{M}}$. Let $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ (resp. $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$) be a lift of $x$ (resp. of $y\in{\mathcal{M}}$). Then $\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $\mu^{-}]\in\Lambda^{-}_{\Gamma}$.
@@
Assuming that $\xi^{+}(\gamma_{-})$ is transverse to $\mu^{-}]$ leads to a contradiction: since $\mu_{k}^{-}\to\mu^{-}\neq 0$, Lemma 7.12 would imply that $\gamma_{k}\mu_{k}^{-}\to 0$, yet $\gamma_{k}\mu_{k}^{-}=e^{t_{k}}f\nu_{k}^{-}$ and $\nu_{k}^{-}\to\nu^{-}\neq 0$. Since $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$, we must have $\mu^{+}]\pitchfork\xi^{-}(\gamma_{-})$, therefore $\gamma_{k}[\mu^{+}_{k}]\to\xi^{+}(\gamma_{+})$ by (A3). This means that $f[\nu^{+}]=\xi^{+}(\gamma_{+})$, and in particular $\nu^{+}]\in\Lambda^{+}_{\Gamma}$.
This implies that $\xi^{+}(\gamma_{+})\pitchfork f[\nu^{-}]$, so $\gamma_{k}^{-1}f[\nu^{-}_{k}]\to\xi^{-}(\gamma_{-})$, i.e. $\mu^{-}]=\xi^{-}(\gamma_{-})\in\Lambda^{-}_{\Gamma}$. ∎
@@
Let $x\in{\mathcal{M}}$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Then $x\in\mathrm{Per}(\phi^{t})$ if and only if there is $\gamma\in\Gamma\setminus\{e\}$ such that $\nu^{+}]=\xi^{+}(\gamma^{+})$ and $\nu^{-}]=\xi^{-}(\gamma^{-})$. [^3] Recall that we assume $\Gamma$ to be torsion-free here, so that any $\gamma\in\Gamma\setminus\{e\}$ has infinite order. In this case, if $\gamma$ is primitive, then the period of $x$ is $\mathcal{J}^{\mathbf{s}}_{\Theta}(\gamma)$.
@@
Assume that $x\in\mathrm{Per}(\phi^{t})$, and let $T>0$ be such that $\phi^{T}(x)=x$. There is $\gamma\in\Gamma\setminus\{e\}$ such that $(e^{-T}\nu^{+},e^{T}\nu^{-})=\gamma\cdot(\nu^{+},\nu^{-})$. If $\nu^{-}]$ were to be transverse to $\xi^{+}(\gamma^{-})$, then Lemma 7.12 would imply that $\gamma^{k}\cdot\nu^{-}\to 0$ as $k\to+\infty$, which is a contradiction with $\gamma^{k}\cdot\nu^{-}=e^{kT}\nu^{-}$. As $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$, we must have $\nu^{+}]\pitchfork\xi^{-}(\gamma^{-})$, so $\gamma^{k}[\nu^{+}]\to\xi^{+}(\gamma^{+})$ as $k\to+\infty$, i.e. $\nu^{+}]=\xi^{+}(\gamma^{+})$. Since $\nu^{-}]\pitchfork[\nu^{+}]=\xi^{+}(\gamma^{+})$, we also find that $\gamma^{k}[\nu^{-}]\to\xi^{-}(\gamma^{-})$ as $k\to$-$\infty$, therefore $\nu^{-}]=\xi^{-}(\gamma^{-})$.
If the assumption is now that $\nu^{+}]=\xi^{+}(\gamma^{+})$ and $\nu^{-}]=\xi^{-}(\gamma^{-})$ for some element $\gamma\in\Gamma\setminus\{e\}$, then combining (6.5), (3.21) and (3.19) yields $\phi^{w_{{\mathbf{s}},\Theta}(\lambda(\gamma))}(x)=x$. ∎
@@
Any flow satisfies $\overline{\mathrm{Per}(\phi^{t})}\subset\mathcal{NW}(\phi^{t})$. Let $x\in\mathcal{NW}(\phi^{t})$, and consider sequences $x_{k}\to x$ in $\mathcal{M}$} and $t_{k}\to+\infty$ in $\mathbb{R}$ such that $\phi^{t_{k}}(x_{k})\to x$. If $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ is a lift of $x$, then Lemma 8.1 shows that $\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$, i.e. $\mathcal{NW}(\phi^{t})\subset\mathcal{K}$.
@@
Let $U\subset{\mathcal{M}}$ be a relatively compact subset containing $\mathcal{K}$, and consider $y\in\overline{\bigcap_{t\in\mathbb{R}}\phi^{t}(U)}$ and a lift $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$. Consider first sequences $t_{k}\to+\infty$ and $x_{k}\in U$ such that $\phi^{t_{k}}(x_{k})\to y$, and assume without loss of generality that $x_{k}$ has a limit $x\in{\mathcal{M}}$. Lemma 8.1 shows that $\mu^{-}]\in\Lambda^{-}_{\Gamma}$. Consider now sequences $s_{k}\to$-$\infty$ and $z_{k}\in U$ such that $y=\lim\phi^{s_{k}}(z_{k})$, and assume that $z_{k}\to z\in{\mathcal{M}}$. Applying Lemma 8.1 to $\phi^{-s_{k}}\left(\phi^{s_{k}}(y)\right)$ shows that $\mu^{+}]\in\Lambda^{+}_{\Gamma}$, hence $y\in\mathcal{K}$.
@@
Since $\mathcal{K}$ is compact and $\phi^{t}$-invariant, we automatically get $\mathcal{K}\subset{\mathcal{T}}(\phi^{t})$. Let $x\in{\mathcal{T}}(\phi^{t})$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Let $y\in{\mathcal{M}}$ be an $\omega$-limit point of $x$, i.e. $y=\lim_{k\to+\infty}\phi^{t_{k}}(x)$ for some sequence $t_{k}\to+\infty$ (it exists because $x$ is trapped). Then Lemma 8.1 shows that $\nu^{+}]\in\Lambda^{+}_{\Gamma}$. The same reasoning applied to an $\alpha$-limit point of $x$ (i.e. some $\lim\phi^{s_{k}}(x)$ with $s_{k}\to$-$\infty$) shows that $\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$.
[…]